# PepCompass: Navigating peptide embedding spaces using Riemannian Geometry

### v4

@@ 2.7 PoGS - Potential-minimizing Geodesic Search
$\mathcal{E}_{\lambda,\mu}(Z)\;=\;\underbrace{\sum_{k=0}^{N-1}\|X_{k+1}-X_{k}\|_{2}^{2}}_{\text{kinetic term: geometric similarity}}\;+\;\lambda\,\underbrace{\sum_{k=0}^{N}\Phi(X_{k})}_{\text{potential term: property bias}}\;+\;\mu\,\underbrace{\sum_{k=0}^{N-1}\|z_{k+1}-z_{k}\|_{2}^{2}}_{\text{latent regularizer}},$ (6)
@@
To perform optimization and search, we initialize $Z$ by straight-line interpolation in latent space, and later optimize $\mathcal{E}_{\lambda,\mu}(Z)$ w.r.t. $Z$ using ADAM solver. During optimization, only the interior points $z_{1},\dots,z_{N-1}$ are updated. Given an optimized path $Z$, we decode each $z_{i}$ to a peptide $\mathtt{p}_{i}$ and remove consecutive duplicates, obtaining a peptide path $\left(\mathtt{p}^{\prime}_{0},\dots,\mathtt{p}^{\prime}_{N^{\prime}}\right)$ of length $N^{\prime}$. We call a peptide $\mathtt{p}^{\prime}_{k}$ a seed if its potential satisfies $\Phi(\mathtt{p}^{\prime}_{k})\leq\theta_{\text{pot}}$ for a threshold $\theta_{\text{pot}}$. A peptide $\mathtt{p}^{\prime}_{k}$ is called a well if it is a seed and also a local minimum of the potential $\Phi$ along the peptide path. The detailed algorithm is presented in Appendix G (Algoritm 6).
@@ 3.1 PoGS evaluation
For each pair $z_{a}$ and $z_{b}$, we compared three paths between $z_{a}$ and $z_{b}$: straight Euclidean interpolation, PoGS without potential, and full PoGS ; Sec. 2.7), measuring chord latent and ambient lengths, decoded peptide path lengths, and property-based counts of seeds and wells. PoGS hyperparametrs and metrics are described in Appendix G. PoGS achieved shorter ambient paths, and substantially more seeds and wells (Table 1), showing that property-aware potentials enrich trajectories for active candidates.
@@ 3.2 LE-BO evaluation
We next evaluated our LE-BO algorithm on a black-box peptide optimization task with a budget of 1400 evaluations. Optimization was initialized from four seed peptides derived using PoGS for $\lambda~=0.01$ and $\mu=0.1$ (Sec. 2.7): KY14, KF16, KK16, FL14, as well as two previously described AMPs (mammuthusin-3, hydrodamin-2[53]. The optimization procedure was repeated 10 times, and for each run the best value achieved across all optimization steps was recorded and reported in Table 2. LE-BO hyperparameters are described in Appendix I.
@@
As the optimized black-box function, we used the APEX MIC regressor [53], with the objective of minimization. We minimize the average $\log_{2}$ MIC across three E. coli strains, reporting the mean over the best results from 10 repeated optimization runs (Table 2).
LE-BO achieved the best performance on five out of six prototypes, outperforming all baselines except for FL14, where it ranked second. The Ablated LE-BO variant with Euclidean walk and enabled mutations achieved the second-best performance on four out of six prototypes, underscoring the importance of mutation enumeration in the optimization process. These ablations further confirmed that both Riemannian random walks and mutations are essential for consistently achieving low MIC values.
3.3 Wet-lab validation
To validate the computational predictions from PoGS and LE-BO, we conducted comprehensive in vitro antimicrobial testing of the generated peptides. A total of 29 novel peptides were experimentally evaluated: 4 seed peptides discovered through PoGS bi-prototype geodesics with property-aware potentials, and 25 analogs derived from these 4 seeds through LE-BO optimization for E. coli activity. These peptides were tested against a panel of 19 bacterial strains, including 8 multidrug-resistant (MDR) isolates (Appendix J), to assess both broad-spectrum activity and efficacy against clinically relevant resistant pathogens. We compared the success rate of PepCompass to previous methods that were also validated experimentally (HydrAMP [50], AMP-Diffusion [51], CLaSS [12], Joker [40]). To this end, for each activity threshold, we computed the fraction of tested peptides that were active against at least one bacterial strain with this activity thershold.
[…]
@@ Conclusions
By leveraging Riemannian latent geometry, interpretable tangent-space mutations, and potential-augmented geodesics, PepCompass enables efficient navigation of and optimization within peptide space across global and local scales. One of possible limitation of our approach is the reliance on the Eucleadian-distance based metric on decoder outputs in the ambient space. However, any other metrics that would better capture the ambient manifold could easily be incorporated. Already now, our computational experiments demonstrate superior performance over state-of-the-art baselines, and wet-lab validation confirms unprecedented success rates, with all tested peptides showing activity in vitro, including activity against multidrug-resistant pathogens. These results establish geometry-aware exploration as a powerful new paradigm for controlled generative design in vast biological spaces.
Reproducibility Statement
All proofs, together with their explanations and underlying assumptions, are provided in Appendices B, C and D. All implementation details and hyperparameters used in the experiments are listed in Appendices E, F, G, H and I. Additionally, we release the source code and experimental configurations necessary to reproduce the key results. Full details of wet-lab validation procedure are described in Appendix J.
Ethics statement
This work includes methods generally applicable to peptide design. An example of potential malicious use of PepCompass would include optimization of peptide toxicity. However, the intention of this paper is to instead provide tools facilitating the design of therapeutic peptides.
LLM usage
Large Language Models (LLMs) were used in this work to improve the clarity and structure of the text. Their use was limited to rephrasing and stylistic refinement. In addition, LLMs were employed to support the search of related work, helping to verify the accuracy of claims about prior research.
Appendix
@@ Appendix A Related work
Deep generative models can be endowed with Riemannian structure [16] by pulling back the ambient metric through the decoder Jacobian [3, 4, 44]. This allows distances and geodesics to reflect data geometry rather than Euclidean latent coordinates. However, most approaches assume a single smooth manifold of fixed dimension. Evidence from both theory and experiments shows that data often lie on unions of manifolds with varying intrinsic dimension or CW-complex structures [36, 9, 56]. In practice, existing methods either restrict the latent to very low dimensions ($\leq 8$) or inflate the metric with variance terms to ensure full rank [4, 15], but these ignore extrinsic geometry. Our model instead decomposes the latent into Riemannian submanifolds of varying dimensions, enabling principled geometry across heterogeneous regions.
@@
Latent Brownian motion has been used as a prior for VAEs [31] and for Riemannian score-based modeling [14]. But these approaches rely on first-order updates. Convergence results for geodesic random walks show that correct Riemannian and sub-Riemannian Brownian motion requires second-order approximations [42, 26]. Our method explicitly incorporates this requirement, yielding diffusion-consistent walks where previous methods diverge.
@@
Tangent-space analysis has mostly been applied in vision, where interpretable latent directions are discovered in GANs or diffusion models via Jacobian or eigen decompositions [45, 60, 57, 1]. Frames induced by augmentations provide another lens on local tangent geometry [34]. We are the first to provide an interpretable tangent space in peptide sequence models, where tangent vectors correspond directly to biologically meaningful mutations.
@@
Geodesics are widely used for interpolation and counterfactual reasoning in latent space [39, 6]. Yet these are typically free geodesics. We extend the concept with potentials, leveraging the Jacobi metric [22] so that peptide traversals account for both geometry and biochemical preferences.
@@
Geometry-aware latent models have been used for chemical-space exploration [59, 58], molecule optimization [20], and protein sequence modeling [10, 15]. Our approach complements these by combining: (i) varying-dimension latent decomposition, (ii) second-order consistent Brownian walks, (iii) interpretable tangent spaces via mutations, and (iv) potential-augmented geodesics tailored to peptide design.
Appendix B $\kappa$-stable dimension of the PepCVAE [13] and HydrAMP [50] models
@@
To quantify how the $\kappa$-stable dimension varies across the latent space, we sampled 1,0001{,}000 points from the HydrAMP [50] training set and computed their $\kappa$-stable dimensions under both the PepCVAE [13] and HydrAMP [50] models. We set $\kappa=10^{-8}$, corresponding to the precision of the float32 format commonly used in neural network computations. This choice ensures that no eigenvalue of the inverse metric tensor exceeds $10^{8}$, thereby avoiding numerical instabilities.
@@ Appendix C Construction of $\kappa$-Stable Manifolds
$\bar{W}^{\kappa}_{z}\cap(W_{z}^{\kappa})^{\perp}\subset\operatorname{span}\{V(z)_{:,k_{z}^{\kappa}},\dots,V(z)_{:,d-1}\}.$
@@
$\leq\kappa\Big(a_{k_{z}^{\kappa}}^{2}\|V(z)_{:,k_{z}^{\kappa}}\|^{2}+\dots+a_{d-1}^{2}\|V(z)_{:,d-1}\|^{2}\Big)$ (12)
$=\kappa,$ (13)
@@ C.1 On the local instability of $\kappa$-stable dimension
This observation prevents a direct application of the Frobenius or constant rank theorem, which require a rank function that is constant in a neighborhood. Our construction in Sec. 2.1 circumvents this issue by working with an open set $W^{\kappa}_{z}$ around $z$ on which the rank remains constant, thereby ensuring that both $M^{\kappa}_{z}$ and $\operatorname{Dec}(W^{\kappa}_{z})$ are smooth $k^{\kappa}_{z}$-dimensional submanifolds.
Appendix D Proof of Theorem 1

### v5

@@ 2.7 PoGS - Potential-minimizing Geodesic Search
$\begin{split}\mathcal{E}_{\lambda,\mu}(Z)\;=\;&\underbrace{\sum_{k=0}^{N-1}\|X_{k+1}-X_{k}\|_{2}^{2}}_{\text{kinetic term: geometric similarity}}\;+\;\lambda\,\underbrace{\sum_{k=0}^{N}\Phi(X_{k})}_{\text{potential term: property}}\\ &\;+\;\mu\,\underbrace{\sum_{k=0}^{N-1}\|z_{k+1}-z_{k}\|_{2}^{2}}_{\text{latent regularizer}},\end{split}$ (6)
@@
To perform optimization and search, we initialize $Z$ by straight-line interpolation in latent space, and later optimize $\mathcal{E}_{\lambda,\mu}(Z)$ w.r.t. $Z$ using ADAM solver. During optimization, only the interior points $z_{1},\dots,z_{N-1}$ are updated. Given an optimized path $Z$, we decode each $z_{i}$ to a peptide $\mathtt{p}_{i}$ and remove consecutive duplicates, obtaining a peptide path $\left(\mathtt{p}^{\prime}_{0},\dots,\mathtt{p}^{\prime}_{N^{\prime}}\right)$ of length $N^{\prime}$. We call a peptide $\mathtt{p}^{\prime}_{k}$ a seed if its potential satisfies $\Phi(\mathtt{p}^{\prime}_{k})\leq\theta_{\text{pot}}$ for a threshold $\theta_{\text{pot}}$. A peptide $\mathtt{p}^{\prime}_{k}$ is called a well if it is a seed and also a local minimum of the potential $\Phi$ along the peptide path. The detailed algorithm is presented in Appendix I (Algoritm 6).
@@ 3.1 PoGS evaluation
For each pair $z_{a}$ and $z_{b}$, we compared three paths between $z_{a}$ and $z_{b}$: straight Euclidean interpolation, PoGS without potential, and full PoGS ; Section 2.7), measuring chord latent and ambient lengths, decoded peptide path lengths, and property-based counts of seeds and wells. PoGS hyperparametrs and metrics are described in Appendix I. PoGS achieved shorter ambient paths, and substantially more seeds and wells (Table 1), showing that property-aware potentials enrich trajectories for active candidates. In Appendix K we present additional analyses that further validate these findings, demonstrating that the observed gains are robust to the choice of ambient metric, set of prototypes and to the potential function.
@@ 3.2 LE-BO evaluation
We next evaluated our LE-BO algorithm on a black-box peptide optimization task with a budget of 1400 evaluations. Optimization was initialized from four seed peptides derived using PoGS for $\lambda\penalty 10000\ =0.01$ and $\mu=0.1$ (parameters described in Section 2.7): KY14, KF16, KK16, FL14, as well as two previously described AMPs (mammuthusin-3, hydrodamin-2) (Wan et al., 2024). As the optimized black-box function, we used the APEX MIC regressor (Wan et al., 2024), with the objective of minimization. We minimize the average $\log_{2}$ MIC across three E. coli strains, reporting the mean over the best results from 10 repeated optimization runs (Table 2). LE-BO hyperparameters are described in Appendix L.
@@
As reported in Table 2, LE-BO achieved the best performance for five out of six prototypes, outperforming all baselines except for FL14, where it ranked second. The Ablated LE-BO variant with Euclidean walk and enabled mutations achieved the second-best performance on four out of six prototypes, underscoring the importance of mutation enumeration in the optimization process. These ablations further confirmed that both Riemannian random walks and mutations are essential for consistently achieving low MIC values. Additional analyses in Appendix O provide a geometric explanation for these trends. By adapting to the local decoder-induced geometry, Riemannian LE-BO identifies substantially more local candidates for Bayesian optimization, particularly in regions with higher $\kappa$-stable dimensions, than Euclidean-based (Figure 8), explaining its stronger optimization performance. Notably, these high-dimensional regions are associated with lower charge and aromaticity (Figure 9), and thus less favorable antimicrobial properties, where richer local enumeration increases the likelihood that such deficiencies can be corrected during optimization.
To further assess the reliability and generality of LE-BO, we conduct an extensive set of ablation and sensitivity analyses, summarized in Appendix N. These experiments evaluate the impact of key algorithmic components, including the stability parameters $\kappa_{\textsc{SORBES}}$ and $\kappa_{\textsc{MUTANG}}$, the removal of the ROBOT component, variation across different initial seeds, and the use of alternative oracle models. Since LE-BO employs only a single-step SORBES-SE walk, the contraction parameter $\alpha$ does not affect the resulting dynamics and is therefore omitted from the ablation study. Together, these analyses demonstrate that the performance gains of LE-BO are stable across algorithmic choices and experimental settings.
@@
3.3 Wet-lab validation
To validate the computational predictions from PoGS and LE-BO, we conducted comprehensive in vitro antimicrobial testing of the generated peptides. A total of 29 novel peptides were experimentally evaluated: 4 seed peptides discovered through PoGS bi-prototype geodesics with property-aware potentials, and 25 analogs derived from these 4 seeds through LE-BO optimization for E. coli activity. These peptides were tested against a panel of 19 bacterial strains, including 8 multidrug-resistant (MDR) isolates (Table 12, Appendix P), to assess both broad-spectrum activity and efficacy against clinically relevant resistant pathogens. We compared the success rate of PepCompass to previous methods that were also validated experimentally (HydrAMP (Szymczak et al., 2023), AMP-Diffusion (Torres et al., 2025), CLaSS (Das et al., 2021), Joker (Porto et al., 2018)). To this end, for each activity threshold, we computed the fraction of tested peptides that were active against at least one bacterial strain with this activity threshold.
[…]
@@ Conclusions
By leveraging Riemannian latent geometry, interpretable tangent mutations, and potential-augmented geodesics, PepCompass enables efficient navigation and optimization within peptide space across global and local scales. One potential limitation of the current implementation is the use of a Euclidean metric in the decoder output space. While seemingly simplistic, it induces a structurally conservative geometry that is mutationally uniform and conservative to a non-local insertions and deletions; see Appendix D for a rationale and potential extensions to more biologically informed metrics, which we plan to explore in future work. Already now, our computational experiments demonstrate superior performance over state-of-the-art baselines, and wet-lab validation confirms unprecedented success rates, with all tested peptides showing in vitro activity, including against multidrug-resistant pathogens. These results establish geometry-aware exploration as a powerful paradigm for controlled generative design in vast biological spaces.
Impact Statement
This work advances the field of machine learning by developing a principled, geometry-aware framework (PepCompass) for navigating high-dimensional generative latent spaces. While the primary technical contribution concerns methods for effective exploration and optimization of peptide design spaces, the application domain: machine-aided antimicrobial peptide discovery, has clear implications for human health and global societal wellbeing.
Antimicrobial resistance (AMR) is a pressing public health challenge worldwide, contributing to increased morbidity, mortality, and healthcare costs. Machine learning that accelerates de novo design of effective antimicrobial agents can contribute positively by expanding the repertoire of candidate therapeutics and potentially reducing time and cost in early drug discovery. By facilitating systematic exploration of peptide design spaces, this work may help identify novel active compounds that would be difficult or costly to find through traditional methods alone.
At the same time, broader impacts must be considered. Techniques that enhance biological sequence design raise questions around dual-use; for example, the same generative design mechanisms could theoretically be repurposed, intentionally or unintentionally, to optimize sequences with harmful biological activity. While our current work focuses on antimicrobial peptides against pathogenic bacteria, any general framework for biological sequence optimization warrants careful consideration of misuse scenarios. Responsible research practice in this area includes transparency in model limitations, rigorous experimental validation, and engagement with domain experts from biology, biosecurity, and ethics to assess risks associated with misuse.
We note that no human data, personally identifiable information, or sensitive attributes are used or generated in this work. The datasets employed are publicly available peptide sequence collections. Our method does not automate experimental execution; in vitro validation requires separate laboratory resources and domain expertise.
@@ Appendix A Related work
Deep generative models can be endowed with Riemannian structure (Flaherty and do Carmo, 2013) by pulling back the ambient metric through the decoder Jacobian (Shao et al., 2017; Arvanitidis et al., 2020, 2021). This allows distances and geodesics to reflect data geometry rather than Euclidean latent coordinates. However, most approaches assume a single smooth manifold of fixed dimension. Evidence from both theory and experiments shows that data often lie on unions of manifolds with varying intrinsic dimension or CW-complex structures (Lou, 2021; Brown et al., 2023; Wang and Wang, 2024). In practice, existing methods either restrict the latent to very low dimensions ($\leq 8$) or inflate the metric with variance terms to ensure full rank (Arvanitidis et al., 2020; Detlefsen et al., 2022), but these ignore extrinsic geometry. Our model instead decomposes the latent into Riemannian submanifolds of varying dimensions, enabling principled geometry across heterogeneous regions.
@@
Latent Brownian motion has been used as a prior for VAEs (Kalatzis et al., 2020) and for Riemannian score-based modeling (De Bortoli et al., 2022). But these approaches rely on first-order updates. Convergence results for geodesic random walks show that correct Riemannian and sub-Riemannian Brownian motion requires second-order approximations (Schwarz et al., 2023; Herrmann et al., 2023). Our method explicitly incorporates this requirement, yielding diffusion-consistent walks where previous methods diverge.
@@
Tangent-space analysis has mostly been applied in vision, where interpretable latent directions are discovered in GANs or diffusion models via Jacobian or eigen decompositions (Shen et al., 2020; Dombrowski et al., 2021; Zhu et al., 2021; Park et al., 2023). Frames induced by augmentations provide another lens on local tangent geometry (Kvinge et al., 2023). We are the first to provide an interpretable tangent space in peptide sequence models, where tangent vectors correspond directly to biologically meaningful mutations.
@@
Geodesics are widely used for interpolation and counterfactual reasoning in latent space (Pegios et al., 2024; Kim et al., 2024). Yet these are typically free geodesics. We extend the concept with potentials, leveraging the Jacobi metric (Gibbons, 2015) so that peptide traversals account for both geometry and biochemical preferences.
@@
Geometry-aware latent models have been used for chemical-space exploration (Du et al., 2022; Wei et al., 2024), and protein sequence modeling (Lys Sanz Moreta, Ola Rønning, Ahmad Salim Al-Sibahi, Jotun Hein, Douglas Theobald, Thomas Hamelryck, 2022; Detlefsen et al., 2022). Our approach complements these by combining: (i) varying-dimension latent decomposition, (ii) second-order consistent Brownian walks, (iii) interpretable tangent spaces via mutations, and (iv) potential-augmented geodesics tailored to peptide design.
Appendix B $\kappa$-stable dimension of the PepCVAE (Das et al., 2018) and HydrAMP (Szymczak et al., 2023) models
@@
To quantify how the $\kappa$-stable dimension varies across the latent space, we sampled 1,0001{,}000 points from the HydrAMP (Szymczak et al., 2023) training set and computed their $\kappa$-stable dimensions under both the PepCVAE (Das et al., 2018) and HydrAMP (Szymczak et al., 2023) models. We set $\kappa=10^{-8}$, corresponding to the precision of the float32 format commonly used in neural network computations. This choice ensures that no eigenvalue of the inverse metric tensor exceeds $10^{8}$, thereby avoiding numerical instabilities.
@@ Appendix C Construction of $\kappa$-Stable Manifolds
$\bar{W}^{\kappa}_{z}\cap(W_{z}^{\kappa})^{\perp}\subset\operatorname{span}\{V(z)_{:,k_{z}^{\kappa}},\dots,V(z)_{:,d}\}.$
@@
$\leq\kappa^{2}\Big(a_{k_{z}^{\kappa}}^{2}\|V(z)_{:,k_{z}^{\kappa}}\|^{2}+\dots+a_{d-1}^{2}\|V(z)_{:,d-1}\|^{2}\Big)$ (12)
$=\kappa^{2},$ (13)
@@ C.1 On the local instability of $\kappa$-stable dimension
This observation prevents a direct application of the Frobenius or constant rank theorem, which require a rank function that is constant in a neighborhood. Our construction in Section 2.1 circumvents this issue by working with an open set $W^{\kappa}_{z}$ around $z$ on which the rank remains constant, thereby ensuring that both $M^{\kappa}_{z}$ and $\operatorname{Dec}(W^{\kappa}_{z})$ are smooth $k^{\kappa}_{z}$-dimensional submanifolds.
Appendix D Mutational Interpretation of the Ambient Euclidean Metric
In this section we elaborate on the interpretation of the Euclidean metric used in the decoder output space $\hat{X}\subset\mathbb{R}^{LA}$ and its implications for the geometry of peptide modifications, as well as on possible metric extensions.
D.1 Euclidean mutation geometry
A discrete peptide sequence $\mathtt{p}$ of length $L$ over an alphabet $\mathcal{A}$ of size $A$ is represented as a vector $\texttt{oh}(\mathtt{p})\in\{0,1\}^{LA}\subset\mathbb{R}^{LA}$ obtained by concatenating $L$ one-hot vectors,
$\texttt{oh}(\mathtt{p})=(e_{a_{1}},\dots,e_{a_{L}}),$
where $e_{a_{\ell}}\in\mathbb{R}^{A}$ denotes the canonical basis vector corresponding to amino acid $a_{\ell}$ at position $\ell$.
Let $\texttt{mut}_{a\rightarrow b,\ell}$ denote the substitution of amino acid $a$ by $b$ at position $\ell$. For any peptide $\mathtt{p}$ with $a_{\ell}=a$,
$\texttt{oh}(\texttt{mut}_{a\rightarrow b,\ell}(\mathtt{p}))-\texttt{oh}(\mathtt{p})=\psi_{a\rightarrow b,\ell},$
where
$\psi_{a\rightarrow b,\ell}=(0,\dots,e_{b}-e_{a},\dots,0)\in\mathbb{R}^{LA}.$
For two substitutions $\texttt{mut}_{a\rightarrow b,\ell_{1}}$ and $\texttt{mut}_{c\rightarrow d,\ell_{2}}$,
$\langle\psi_{a\rightarrow b,\ell_{1}},\psi_{c\rightarrow d,\ell_{2}}\rangle=\mathbf{1}_{\{\ell_{1}=\ell_{2}\}}\big(\mathbf{1}_{\{a=c\}}+\mathbf{1}_{\{b=d\}}-\mathbf{1}_{\{a=d\}}-\mathbf{1}_{\{b=c\}}\big).$
Consequently,
$\|\psi_{a\rightarrow b,\ell}\|_{2}^{2}=2,$
substitutions acting on different positions are orthogonal, and substitutions at the same position have constant pairwise cosine similarity $1/2$ (corresponding to an angle of $30^{\circ}$). Thus, the Euclidean metric induces a uniform and isotropic geometry over substitutional mutations.
This implies that using a Euclidean metric in the ambient amino-acid probability space can be interpreted as imposing a non-informative prior over mutation directions.
D.2 Euclidean insertion and deletion geometry
Let $\texttt{ins}_{a,\ell}$ and $\texttt{del}_{\ell}$ denote insertion of amino acid $a$ at position $\ell$ and deletion at position $\ell$, respectively. Terminal insertions and deletions can be treated as substitutions involving a padding symbol pad and therefore have squared norm $2$, and are orthogonal to other substitutions.
For internal insertions and deletions, shifting of downstream residues induces a perturbation, whose squared Euclidean norm satisfies
$\|\texttt{oh}(\texttt{ins}_{a,\ell}(\mathtt{p}))-\texttt{oh}(\mathtt{p})\|_{2}^{2}=2\Big(\mathbf{1}_{\{\mathtt{p}_{\ell}=a\}}+\sum_{i=\ell}^{L-1}\mathbf{1}_{\{\mathtt{p}_{i+1}\neq\mathtt{p}_{i}\}}\Big)\leq 2(L-\ell+1),$
and
$\|\texttt{oh}(\texttt{del}_{\ell}(\mathtt{p}))-\texttt{oh}(\mathtt{p})\|_{2}^{2}=2\Big(1+\sum_{i=\ell}^{L-1}\mathbf{1}_{\{\mathtt{p}_{i+1}\neq\mathtt{p}_{i}\}}\Big)\leq 2(L-\ell+1).$
The potentially large displacement reflects the non-local nature of internal insertions and deletions, and appropriately penalizes mutations which are likely to induce global structural rearrangements, such as register shifts or changes in secondary structure.
D.3 Extension to biologically meaningful ambient metrics by mutation kernels
Let
$\texttt{Mut}=\{\texttt{mut}_{a\rightarrow b,\ell}:a,b\in\mathcal{A},\ \ell\in\{1,\dots,L\}\}$
be the set of substitution operators, and consider a positive-definite mutation kernel
$\mathcal{K}_{\texttt{Mut}}:\texttt{Mut}\times\texttt{Mut}\rightarrow\mathbb{R}.$
Here, diagonal terms of $\mathcal{K}_{\texttt{Mut}}$ control mutation magnitudes, while off-diagonal terms encode similarity between distinct mutations, enabling incorporation of biochemical substitution preferences and positional coupling into the ambient geometry.
Define
$V_{\texttt{Mut}}=\operatorname{span}\{\psi_{a\rightarrow b,\ell}:\texttt{mut}_{a\rightarrow b,\ell}\in\texttt{Mut}\}.$
For $v_{1}=\sum_{i}c_{i}\psi_{i}$ and $v_{2}=\sum_{j}d_{j}\psi_{j}$, define
$\mathcal{K}_{\mathcal{K}_{\texttt{Mut}}}^{V_{\texttt{Mut}}}(v_{1},v_{2})=\sum_{i,j}c_{i}d_{j}\,\mathcal{K}_{\texttt{Mut}}(\texttt{mut}_{i},\texttt{mut}_{j}).$
Every $w\in\mathbb{R}^{LA}$ decomposes uniquely as $w=w^{V_{\texttt{Mut}}}+w^{V_{\texttt{Mut}}^{\perp}}$ under the standard Euclidean inner product. Define the ambient kernel
$\mathcal{K}_{\mathcal{K}_{\texttt{Mut}}}^{\mathbb{R}^{LA}}(w_{1},w_{2})=\mathcal{K}_{\mathcal{K}_{\texttt{Mut}}}^{V_{\texttt{Mut}}}(w_{1}^{V_{\texttt{Mut}}},w_{2}^{V_{\texttt{Mut}}})+\langle w_{1}^{V_{\texttt{Mut}}^{\perp}},w_{2}^{V_{\texttt{Mut}}^{\perp}}\rangle.$
This construction transfers biologically meaningful similarity between mutations to a Riemannian metric on the ambient space. We plan to explore such metrics in future work.
Extension of $\kappa$-stable manifold construction to kernel-based ambient metrics
Let $e_{i}\in\mathbb{R}^{LA}$, $i=1,\dots,LA$, denote the canonical basis vectors, and define the Gram matrix
$G_{\mathcal{K}}[i,j]=\mathcal{K}_{\mathcal{K}_{\texttt{Mut}}}^{\mathbb{R}^{LA}}(e_{i},e_{j}).$
Let
$G_{\mathcal{K}}=U_{\mathcal{K}}\Sigma_{\mathcal{K}}U_{\mathcal{K}}^{\top}$
be its eigen decomposition. Define a modified decoder
$\hat{\operatorname{Dec}}_{\mathcal{K}}(z)=\Sigma_{\mathcal{K}}^{1/2}U_{\mathcal{K}}^{\top}\hat{\operatorname{Dec}}(z).$
Then
$J_{\hat{\operatorname{Dec}}_{\mathcal{K}}}(z)=\Sigma_{\mathcal{K}}^{1/2}U_{\mathcal{K}}^{\top}J_{\hat{\operatorname{Dec}}}(z),$
and the induced pullback metric satisfies
$J_{\hat{\operatorname{Dec}}_{\mathcal{K}}}(z)^{\top}J_{\hat{\operatorname{Dec}}_{\mathcal{K}}}(z)=J_{\hat{\operatorname{Dec}}}(z)^{\top}G_{\mathcal{K}}J_{\hat{\operatorname{Dec}}}(z).$
Thus, equipping the ambient space with the Mahalanobis inner product induced by $G_{\mathcal{K}}$ and pulling it back through $\hat{\operatorname{Dec}}$ is equivalent to equipping the ambient space with the standard Euclidean inner product and pulling it back through $\hat{\operatorname{Dec}}_{\mathcal{K}}$ (Flaherty and do Carmo, 2013). In other words, both models induce the same Riemannian metric on latent space. Consequently, the $\kappa$-stable manifold construction from Subsection 2.2 can be applied for $\hat{\operatorname{Dec}}_{\mathcal{K}}$, and the same argument extends to any fixed Mahalanobis-type ambient metric on $\mathbb{R}^{LA}$.
Appendix E Proof of Theorem 2.1